OpenStudy (anonymous):

is under root 4-y^2 a function of y?

OpenStudy (lgbasallote):

i've been waiting to say this for a looooong time "go back to the physics section @Shane_B :P" lol

OpenStudy (shane_b):

Yea...I just noticed this one is not actually a function to begin with!

OpenStudy (lgbasallote):

not really. i just wanted to say it

OpenStudy (shane_b):

@annie1111: does this look like a function to you?

OpenStudy (lgbasallote):

i dont really know anything about this question. we havent discussed i in class yet :S

OpenStudy (shane_b):

This one is simple...it doesn't pass the vertical line test.

OpenStudy (anonymous):

@annie1111 your question is this ? \[\Large x=\sqrt{4-y^2}\]

OpenStudy (anonymous):

if its then it is the equation of circle and ,it is not function.

OpenStudy (lgbasallote):

hypothetical question @sami-21 do you believe in the mythical f(y)?

OpenStudy (shane_b):

A semicircle anyway

OpenStudy (anonymous):

lets make it full circle ! take square of both sides \[\Large x^2+y^2=4\]

OpenStudy (lgbasallote):

...you never answered my question...

OpenStudy (anonymous):

@lgbasallote its just variable.

OpenStudy (lgbasallote):

i have heard of legendary functions such as f(t), f(a), f(b), f(c), etc.

OpenStudy (lgbasallote):

alhough...i cant say i disagree that \(x = \sqrt{4-y^2}\) is not a function of x

OpenStudy (shane_b):

I just realized that you never see f(z)....but you do see f(x), f(y)...poor, poor z. Always left out.

OpenStudy (lgbasallote):

another interesting folklore is sami's legendary typing speed

OpenStudy (anonymous):

i think he is in Algebra course.wants to know it is function or not. they use vertical line test and the following graph fail vertical line test,s mentioned early by shane.

OpenStudy (lgbasallote):

interestingly the question particularly asked function of y....

OpenStudy (lgbasallote):

but heh what do i know. like i said. i havent learned this stuff in class yet

OpenStudy (anonymous):

it was supposed to be learned in Algebra course.

OpenStudy (shane_b):

@lgbasallote: didn't you post some DiffEq questions before?

OpenStudy (shane_b):

if so, you're taking your classes WAY out of order :P

OpenStudy (anonymous):

yeah if i recall, he was playing with some Laplace stuff.

OpenStudy (lgbasallote):

im an advanced reader

OpenStudy (shane_b):

lol

OpenStudy (lgbasallote):

im a 44 year old dyslexic

OpenStudy (shane_b):

I hated DiffEqs...had a HORRIBLE teacher and although I got an A...that's only because everything was take home. I doubt I learned much of anything :/

OpenStudy (anonymous):

yeah i remember your Professor's Laplace equation :)

OpenStudy (anonymous):

Parth is Advance reader.

OpenStudy (shane_b):

I ride the short bus

OpenStudy (lgbasallote):

parth is a pretentious wannabe lol

OpenStudy (lgbasallote):

seems this still isnt solved huh